# Adaptive Extraction Network for Multivariate Long Sequence Time-Series Forecasting

### v1

@@ Abstract
Models employing CNN architecture have made significant progress in multivariate long sequence time-series forecasting (MLSTF), particularly in modeling local time series characteristics. However, during the MLSTF process, extracting the global time series patterns and understanding the correlations among different variables are highly significant. To address this challenge, we introduce multi-resolution convolution and deformable convolution operations. By enlarging the receptive field using convolution kernels with different dilation factors to capture temporal correlation information across different resolutions, and adaptively adjusting the sampling positions through additional offset vectors, we enhance the network’s ability to capture correlated features between variables. Building upon this, we propose ATVCNet, an adaptive temporal-variable convolutional network designed to effectively model the local/global temporal dependencies and inter-variable dependencies of multivariate time series. Specifically, extracting and fusing time series features at different resolutions, captures both local contextual information and global patterns in the time series. The designed inter-variable feature adaptive extraction module captures the correlation among different variables in the time series. We evaluated the performance of ATVCNet across eight real-world datasets. The results indicate that ATVCNet achieved a performance improvement of approximately 63.4% over state-of-the-art MLSTF models.
@@ 1 Introduction
Multivariate long sequence time-series forecasting (MLSTF) stands as one of the most critical challenges in time series analysis, finding extensive applications in fields such as energy [1, 2, 3], data center infrastructure [4], traffic flow [5, 6], finance [7], and more. Its core objective lies in leveraging past time series data to forecast the changing trends of multiple correlated time series over a future time horizon. In recent years, research focus on forecasting methods for multivariate time series has gradually shifted towards Transformer-based models with multi-encoder-decoder architectures and multi-head self-attention mechanisms. This is because Transformer models exhibit strong capabilities in capturing temporal dependencies among long sequences. Consequently, a series of Transformer-based model-driven approaches have emerged, including PatchTST [7], iTransformer [8], and Crossformer [9]. Although Transformer-based models perform well in MLSTF due to their ability to extract temporal information, DLinear [10] and the experiment in Section 4.6 demonstrate that Transformer-based models are often ineffective in capturing global temporal correlations in time series. Additionally, the utilization of self-attention mechanisms at the bottom levels of Transformer-based models for feature extraction leads to high computational costs.
@@
The latest trend in MLSTF research has begun to challenge the efficiency issues of Transformer-based models in prediction tasks. Models like LightTS [11] and DLinear [10], which employ simple linear structures, have shown superior performance in MLSTF compared to the majority of Transformer-based models. However, these models focus on extracting global correlation information in the horizontal time domain, neglecting the understanding of local contextual information and information between different variables, which is crucial for MLSTF. Although TSMixer [12] addresses the issue of extracting cross-variable information by feature dimension transformation and linear mapping, its utilization of fully connected layers for extracting such information leads to information redundancy.
@@
Recently, prediction models based on convolutional modules have shown outstanding performance in MLSTF [13, 14]. These models typically utilize convolutional layers to extract local features from the time series, which to some extent preserve the structural information of the time series. The commonly used convolutional methods for extracting local features include standard convolution and dilated convolution (Figure 1 a,b). The receptive field of standard convolution usually covers only a fixed-size region within the input sequence, exhibiting regularity and fixedness [15], which leads to relatively low variable dependency between regions formed by other sequences. Dilated convolution [14] extends the standard convolution by introducing dilated factor in the convolution kernel, allowing the kernel to span larger receptive fields to capture features at multiple scales. However, dilated convolutions still possess symmetric receptive fields [16], limiting the ability of convolutional layers to perceive features between asymmetric variables, thus presenting certain limitations in capturing sequence dynamics.
@@
To address the aforementioned challenges, we extract relevant features of the time series from both temporal and variable dimensions. Specifically, for capturing the temporal domain information of the time series, we consider extracting local features and patterns at different resolutions within the time series. To extract features between different variables in the time series, we introduce deformable convolutions [17] (Figure 1 c), commonly used in the field of image processing. However, directly applying deformable convolutions to time series prediction tasks may lead to issues of data mismatch and lack of effective learning of temporal intervals.
@@
Based on the analysis above, we designed ATVCNet, an adaptive temporal-variable convolutional network, for accurately MLSTF. ATVCNet consists of two key modules: the temporal feature extraction module and the inter-variable feature adaptive extraction module. The temporal feature extraction module extracts local contextual information and patterns at different resolutions by using convolutional kernels with different dilation factors, and ensures the position invariance of the time series through aggregation operations and uses adaptive average pooling layers to establish global pattern information. The inter-variable feature adaptive extraction module utilizes deformable convolutions to adaptively adjust the shape of the convolutional kernel based on the sequence features, extracting correlations between variables in different feature directions and enhancing the model’s perception of cross-variable correlations. The main contributions are as follows:
  • We propose ATVCNet, an adaptive temporal-variable convolutional network. Experimental studies on 8 real-world datasets demonstrate that the proposed model outperforms the state-of-the-art (SOTA) ConvTimeNet model by 63.4% in terms of MLSTF.
@@
  • To capture the dependencies among different dimensions, we introduced an inter-variable feature adaptive extraction module. This module employs deformable convolution, which enhances the sampling positions between variables by introducing additional offsets. This allows the network to adaptively adjust the shape of the convolution kernels based on specific data characteristics, thereby capturing features between multivariate data more accurately and significantly reducing redundant features.
The remaining structure of this paper is as follows: In Section 2, we introduce existing methods for time series forecasting and the application of deformable convolutions. Section 3 provides a detailed description of the ATVCNet architecture. Section 4 presents experimental results validating the effectiveness and efficiency of ATVCNet. Finally, in Section 5, we provide a comprehensive summary of the predictive research on the ATVCNet model.
@@ 2 Related work
2.1 Time modeling in deep time series forecasting
Deep models in time series forecasting can generally be categorized into four paradigms: those based on CNNs, RNNs, Transformers, and MLPs. Models based on CNNs use convolutional kernels along the time dimension to capture local temporal patterns, such as ConvTimeNet [13], MICN [14], SCINet [18], and TCN [19]. On the other hand, models based on RNNs utilize recurrent structures to maintain state propagation, as seen in LSSL [20], LSTNet [21], LSTM [22], etc. However, both of these paradigms suffer from the problem of limited receptive fields, which hampers their ability for long-term forecasting. Recently, Transformer-based models have garnered widespread attention due to their capability in capturing long-range dependencies through attention mechanisms. For instance, models like Autoformer [23], PatchTST [7], ETSformer [24], FEDformer [25], Pathformer [26], and iTransformer [8] adopt the idea of time series decomposition, breaking down time series data into simpler sequential data to handle complex time patterns. SMARTformer [27], Pyraformer [28], and CrossFormer [9] propose new attention mechanisms to enhance prediction accuracy. However, as the forecasting horizon increases, the attention mechanism’s extraction capability may diminish. Models based on MLPs utilize simple linear models to extract abstract representations of time series, such as LightTS [11], DLinear [10], TSMixer [12], MLP4Rec [29], FreTS [30] and TimeMixer [31]. However, when dealing with high-dimensional data, the expressive power of linear networks is limited, making them unsuitable for the prediction of multivariate data.
In response to the aforementioned challenge, we developed an efficient module for extracting temporal correlations, which models both global and local aspects.
@@ 2.2 Application of deformable convolutions
Deformable convolutions have found widespread application in the field of image processing. In recent research, there has been a growing trend of combining deformable convolutions with attention mechanisms. For instance, in [32, 33], researchers utilized deformable convolutions to capture more important offset information in specific spatial structures, thus providing more precise recognition capabilities. In [34], due to the presence of complex geometric transformations and feature blurring characteristics in the data, researchers employed A2-DCNet modules to capture remote spatial context information from a global perspective. In [35], researchers demonstrated that attention blocks guided by deformable convolutions could acquire semantic information about spatial positions. Additionally, research utilizing deformable convolutions in visual tasks has yielded promising results [36].
These research findings inspire our exploration of deformable convolution patterns in MLSTF. By introducing deformable convolutions into the MLSTF field, we aim is to more effectively capture the correlation information among different variables in time series data, thereby enhancing prediction accuracy and performance.
@@ 3 Methodology
In this section, we will provide a detailed overview of the ATVCNet architecture. As depicted in Figure 2, the ATVCNet model is designed to effectively extract the time-domain pattern characteristics of time series and complex correlation information between different dimensions from multi-dimensional historical data to achieve accurate long-term prediction of time series. The prediction process of the ATVCNet model mainly consists of the following three stages:
@@
(2) Correlation acquisition: It is crucial for accurate MLSTF prediction to adequately extract and utilize both the horizontal time-domain correlation information of the time series and the correlation information among different variables. We employ a temporal feature extraction module and an inter-variable feature adaptive extraction module to extract explicitly correlated features hidden in the original samples, which can significantly enhance prediction accuracy.
(3) Dynamic forecasting: As time progresses, the distribution of time series data in real-world scenarios is likely to change, which may lead to decreased predictive accuracy of the model. Therefore, to maintain the model’s performance, it is necessary to recalibrate the model parameters when there is a decline in predictive accuracy. This adjustment helps adapt to the features of new samples, ensuring the model’s effectiveness in dynamic forecasting tasks.
Figure 2: Framework of the ATVCNet.
@@ 3.2 Data processing
3.3 Temporal-Variable feature extraction
To comprehensively and effectively extract features from time series, we designed the temporal-variable feature extraction module, which consists of two components: the temporal feature extraction module and the inter-variable feature adaptive extraction module, as shown in Figure 3 (a). The former is utilized to capture temporal dependency information of local and global patterns within the time series (as shown in the $TemporalFeatureExtractionModule$ in Figure 3 (a)), while the latter is employed for effectively extracting correlation information between different variables at various resolutions in the time series (as shown in the $Inter-VariableFeatureAdaptiveExtractionModule$ in Figure 3 (a)).
(a) Correlation Acquisition module.
(b) Inter-Variable Feature Adaptive Extraction Module.
Figure 3: Temporal-Variable feature extraction module.
Temporal Feature Extraction Module Inspired by the concept of SPP [38], the temporal feature extraction module employs one-dimensional (1D) convolutions with varying dilation factors to capture features and patterns at different resolutions within the input multivariate time series data. This approach equips the model with the capability to discern local feature variations: with small dilation factors, the model convolves within a small receptive field, facilitating more precise identification of local patterns and structures, thus aiding in enhancing the network’s comprehension of complex patterns. Conversely, with large dilation factors, the model employs larger convolutional kernels to capture features within a broader receptive field, thereby facilitating the identification of local contextual relationships within the input sequence. Additionally, by aggregating feature and pattern information extracted at different resolutions at the same location, it ensures positional invariance of the data during the processing. To encapsulate global trends and patterns, our model incorporates adaptive average pooling operations. This technique is instrumental in distilling the overarching trends and patterns from the time series data, thus enriching the model’s comprehensive grasp of the temporal dynamics presented.
@@ 3.3 Temporal-Variable feature extraction
Inter-Variable Feature Adaptive Extraction Module To effectively extract the correlation and dynamic relationships between different variables under different resolution features of time series, we introduce deformable convolution. By adding additional learnable offsets to standard convolution, it enables the convolutional kernel to adapt to irregular variations in input features, making it particularly suitable for handling complex dynamic relationships between different variables, as illustrated in Figure 3 (b).
Specifically, we stack the extracted multi-resolution feature maps and the output maps from adaptive average pooling into a tensor $X_{sd}$, which serves as the input to the inter-variable feature adaptive extraction module. We employ deformable convolutions to learn the information between different variables and position offset information in feature graphs with different resolutions, focusing attention on regions with richer information through twist sampling networks. This operation enables the model to more accurately capture and understand dependencies between different dimensions while removing redundant feature information. The mathematical expression of the inter-variable feature adaptive extraction module is as follows:
@@
where $DefConv(p_{0})$ is the value of the output feature map at location $p_{0}$. $x(\cdot)$ represents the input feature map. $\mathcal{W}(p_{n})$ denotes the weights of the convolutional kernel. $\triangle p_{n}$ is the learned offset for each $p_{n}$, differing for each point, allowing the convolutional kernel to adaptively adjust its sampling positions on the input feature map. For more detailed theoretical explanations regarding deformable convolutions, please refer to reference [17]. The inter-variable feature adaptive extraction is described by the following equations:
$\begin{split}&X_{sd}=Stack(\chi_{dilated},\chi_{avg\_pool}),\\ &X_{dc}=DefConv(X_{sd}).\end{split}$ (10)

### v2

# Adaptive Convolutional Forecasting Network Based on Time Series Feature-Driven

@@ Abstract
Time series data in real-world scenarios contain a substantial amount of nonlinear information, which significantly interferes with the training process of models, leading to decreased prediction performance. Therefore, during the time series forecasting process, extracting the local and global time series patterns and understanding the potential nonlinear features among different time observations are highly significant. To address this challenge, we introduce multi-resolution convolution and deformable convolution operations. By enlarging the receptive field using convolution kernels with different dilation factors to capture temporal correlation information at different resolutions, and adaptively adjusting the sampling positions through additional offset vectors, we enhance the network’s ability to capture potential nonlinear features among time observations. Building upon this, we propose ACNet, an adaptive convolutional network designed to effectively model the local and global temporal dependencies and the nonlinear features between observations in multivariate time series. Specifically, by extracting and fusing time series features at different resolutions, we capture both local contextual information and global patterns in the time series. The designed nonlinear feature adaptive extraction module captures the nonlinear features among different time observations in the time series. We evaluated the performance of ACNet across twelve real-world datasets. The results indicate that ACNet consistently achieves state-of-the-art performance in both short-term and long-term forecasting tasks with favorable runtime efficiency.
@@ 1 Introduction
Time-series forecasting (TSF) is one of the most critical challenges in time series analysis, with wide-ranging applications in fields such as energy [1, 2, 3], traffic [4], weather [5], disease [6], and more. Its core objective is to leverage past time series data to forecast changing trends over a future time horizon. Effective feature extraction from time series data is essential for improving forecasting accuracy. However, time series data in real-world scenarios often exhibit strong nonlinear features due to factors such as environmental conditions and equipment usage. This poses significant challenges for effective feature extraction. Firstly, there is an increase in noise and uncertainty: data collected by sensors are highly sensitive to environmental changes and external disturbances, leading to increased noise and uncertainty in the time series. Secondly, the extraction of complex dependencies is difficult: time series data in real-world scenarios may exhibit complex nonlinear and dynamic patterns, and the inherent dependency structures and patterns between time observations cannot be accurately captured and described by simple linear models. Additionally, the distribution of time series data in real-world scenarios changes over time, posing higher demands on local feature extraction to effectively capture short-term dynamic behaviors.
To further illustrate the complex nonlinear characteristics of time series data in real-world scenarios, we performed phase space reconstruction on real datasets from different fields. Figure 1 shows the phase space reconstruction diagrams of these datasets, from which the following points can be observed: 1) the trajectory shapes are complex and variable; 2) the trajectories cluster in multiple specific regions; 3) the trajectories are close to each other and intricately intertwined. These characteristics indicate that there are strong nonlinear relationships and dynamic behaviors between different time observations in the time series, which greatly limit the effective extraction of time series data features.
(a) Raw Time Series (PEMS03).
(b) Raw Time Series (PEMS04).
(c) Raw Time Series (ETTh2).
(d) Raw Time Series (Exchange).
(e) Phase Space (PEMS03).
(f) Phase Space (PEMS04).
(g) Phase Space (ETTh2).
(h) Phase Space (Exchange).
Figure 1: Nonlinear Features of Data.
@@
To effectively extract complex features and patterns in multivariate time series, research focus on multivariate time series forecasting methods has gradually shifted towards Transformer-based models, which employ multi-encoder-decoder architectures and multi-head self-attention mechanisms. This shift is attributed to the strong capabilities of Transformer models in capturing temporal dependencies in long sequences. Consequently, a series of Transformer-based model-driven approaches have emerged, including PatchTST [7], iTransformer [8], and Crossformer [9]. Thanks to the ability of the self-attention mechanism to extract long-term dependencies in time series, transformer-based models perform well in TSF. However, they often neglect local contextual information. Additionally, DLinear [10] and the experiments in Section 4.6 show that transformer-based models are generally ineffective at capturing global temporal correlations in time series. Moreover, using the self-attention mechanism for feature extraction at the core of transformer-based models leads to high computational costs.
@@
The latest trend in TSF research has begun to challenge the efficiency issues of Transformer-based models in prediction tasks. Models like TSMixer [11], LightTS [12] and DLinear [10], which employ simple linear structures, have shown superior performance in TSF compared to the majority of Transformer-based models. However, these models focus on extracting global correlation information in the horizontal time domain, neglecting the understanding of local contextual information and nonlinear information between different time observations, which is crucial for TSF.
@@
Recently, prediction models based on convolutional modules have shown outstanding performance in TSF [13, 14]. These models typically utilize convolutional layers to extract local nonlinear features from the time series, which to some extent preserve the dynamic information of the time series. Common methods used for extracting local features include standard convolution and dilated convolution (Figure 2 a, b). The receptive field of standard convolution typically covers only a fixed-size region within the input sequence, exhibiting regularity and fixedness [15]. Dilated convolution [14] extends standard convolution by introducing a dilation factor in the convolution kernel, allowing the kernel to span larger receptive fields to capture multi-scale features. However, they possess symmetric receptive fields [16], which limits the ability of convolutional layers to perceive features between asymmetric variables, thus presenting certain limitations in capturing nonlinear characteristics and dynamics of sequences.
@@
To address the aforementioned challenges, we focus on the effective extraction of time series features. Specifically, to capture the temporal domain information of the time series, we consider extracting local features and patterns at different resolutions within the time series. To extract nonlinear features between different time observations in the time series, we improve deformable convolutions [17] (Figure 2 c), which are commonly used in the field of image processing. Notably, directly applying deformable convolutions to time series prediction tasks may lead to issues of data mismatch and ineffective learning of temporal intervals.
@@
In this work, we propose ACNet, a framework with a multi-resolution mixing architecture capable of extracting local and global information from complex time series data through a temporal feature extraction module (TFE), and then extracting nonlinear features and patterns through a nonlinear feature adaptive extraction module (NFAE). Finally, it predicts future sequences rapidly using a feedforward network and the pseudo-inverse algorithm. With our meticulously designed architecture, ACNet consistently demonstrates leading performance across all experiments, showing superior efficiency in both long-term and short-term forecasting tasks, and covering a wide range of well-established benchmarks. The main contributions are as follows:
  • We propose ACNet, an adaptive convolutional network for long-term and short-term time series forecasting. Experimental studies on fifteen real-world datasets demonstrate that the proposed model outperforms the state-of-the-art ConvTimeNet/Crossformer models by 63.4% and 12.5% in long-term and short-term time series forecasting, respectively.
@@
  • To extract nonlinear features and patterns from complex time series data, we designed a nonlinear feature adaptive extraction module. This module employs improved deformable convolutions, which adaptively adjust the shape of the convolutional kernel based on the input time series features, enhancing the model’s ability to capture nonlinear characteristics and complex patterns between variables in complex time series, thereby improving the accuracy of time series predictions. Additionally, deformable convolutions allow the network to adaptively modify the convolutional kernel shape according to specific data features, accurately capturing the nonlinear features and patterns between different time observations in the time series, and significantly reducing the extraction of redundant features.
  • To improve the computational efficiency of the model, we combine a single hidden layer feedforward neural network (SLFN) with the pseudo-inverse algorithm, which allows for the rapid computation of the optimal solution, significantly reducing training time and computational resource consumption. This approach ensures that our model maintains high prediction accuracy while achieving greater computational efficiency.
The remaining structure of this paper is as follows: In Section 2, we introduce existing methods for time series forecasting and the application of deformable convolutions. Section 3 provides a detailed description of the ACNet architecture. Section 4 presents experimental results validating the effectiveness and efficiency of ACNet. Finally, in Section 5, we provide a comprehensive summary of the predictive research on the ACNet model.
@@ 2 Related work
2.1 Time modeling in time series forecasting
Deep learning models have achieved significant success in time series forecasting, and they can be primarily categorized into three paradigms: CNN-based models, Transformer-based models, and MLP-based models.
CNN-based models utilize convolutional kernels along the temporal dimension to effectively capture local temporal patterns. For instance, TCN [18] uses one-dimensional dilated convolutions to expand the receptive field and performs well in both short-term and long-term forecasting. MICN [14] employs multi-scale convolution to extract both global and local contextual relationships. SCINet [19] decomposes time series into multiple resolutions dynamically through SCI-Blocks in a recursive downsampling fashion to extract features at multiple scales. TSLANet [20] learns long-term and short-term relationships in the data through convolutional operations. ConvTimeNet [13] adaptively segments time series into patches and integrates deepwise convolution and pointwise convolution operations to capture global sequence dependencies and cross-variable interactions. However, the aforementioned models are limited to extracting nonlinear features within a fixed receptive field, which hinders their ability to capture nonlinear features among asymmetric variables.
Transformer-based models have garnered widespread attention due to their capability in capturing long-range dependencies through attention mechanisms. For instance, models such as Autoformer [21], ETSformer [22] and FEDformer [23] decompose time series data to extract complex patterns. CrossFormer [9] captures temporal and cross-dimensional dependencies in multivariate time series data by first transforming it into a 2D array and then employing two-stage attention layers. PatchTST [7] employs a channel independence strategy specifically designed to extract correlations of each channel in multivariate data. Pathformer [24] divides the time series into patches of different scales and designs attention mechanisms within and between patches. iTransformer [8] effectively captures multivariate correlations by independently embedding each variable in the time series as variable sub-tokens. However, as the forecasting horizon increases, the attention mechanism’s extraction capability may diminish.
Models based on MLPs utilize simple linear models to extract abstract representations of time series, such as LightTS [12], DLinear [10] decomposes time series and utilizes linear network layers for modeling to achieve prediction. TSMixer [11] combines Patch and MLP, enabling it to extract both temporal correlations and inter-channel correlations. FreTS [25] and TimeMixer [26] conduct long-term and short-term forecasting by decoupling historical information of complex patterns in multiscale time series. However, when confronted with high-dimensional data, the expressive capacity of linear networks is constrained, thereby impeding their ability to accurately capture nonlinear features within complex and noisy time series.
@@ 2.2 Application of deformable convolutions
Deformable convolutions have found widespread application in the field of image recognition. Recent research has increasingly combined deformable convolutions with attention mechanisms. For instance, in [27, 28], researchers utilized deformable convolutions to capture significant offset information in specific spatial structures, thereby enhancing recognition precision. In [29], due to the complex geometric transformations and feature blurring present in the data, A2-DCNet modules were employed to capture remote spatial context information from a global perspective. Additionally, [30] demonstrated that attention blocks guided by deformable convolutions could acquire semantic information about spatial positions.
Furthermore, research utilizing deformable convolutions in object detection tasks has yielded promising results [31]. In [32], a background-guided deformable convolutional autoencoder network was proposed, effectively separating anomalies from complex backgrounds and enhancing anomaly detection capabilities. Moreover, [33] showed that stacking deformable convolutions and integrating semantic segmentation improved the understanding of contextual relationships. Finally, [34] illustrated that deformable convolutions, akin to spatial attention, could be employed for multiscale feature extraction, while channel attention was used to identify significant features.
To address the aforementioned challenges, we developed a temporal feature extraction module aimed at capturing local contextual information and global patterns within time series data. Additionally, we were inspired by research on the application of deformable convolutions in other domains, prompting us to explore the possibility of introducing them into the field of TSF. By incorporating deformable convolutions, we are able to adaptively capture nonlinear features and patterns within time series data, thereby improving prediction accuracy and performance.
@@ 3 Methodology
In this section, we will provide a detailed overview of the ACNet architecture. As depicted in Figure 3, the ACNet model is designed to effectively extract the time-domain pattern characteristics of time series and complex nonlinear information between different observations from multi-dimensional historical data to achieve accurate prediction of time series. The prediction process of the ACNet model mainly consists of the following three stages:
@@
(2) Feature acquisition: Fully extracting and utilizing the correlation information and nonlinear features between different time observations in a time series is key to accurately predicting TSF. We employ a temporal feature extraction module and a nonlinear feature adaptive extraction module to extract explicit correlated features hidden in the raw samples, which can significantly enhance prediction accuracy.
(3) Dynamic prediction: As time progresses, the distribution of time series data in real-world scenarios is likely to change, which may lead to decreased predictive accuracy of the model. Therefore, to maintain the model’s performance, it is necessary to recalibrate the model parameters when there is a decline in predictive accuracy. This adjustment helps adapt to the features of new samples, ensuring the model’s effectiveness in dynamic forecasting tasks.
Figure 3: Framework of the ACNet.
@@ 3.2 Data processing
3.3 Feature acquisition
To comprehensively and effectively extract features from time series, we designed the feature acquisition module, which consists of two components: the temporal feature extraction module and the nonlinear feature adaptive extraction module, as shown in Figure 4 (a). The former is utilized to capture temporal dependency information of local and global patterns within the time series (as shown in the $TemporalFeatureExtractionModule$ in Figure 4 (a)), while the latter is employed to effectively extract nonlinear information between different time observations and correlations among different variables at various resolutions in the time series. (as shown in the $NonlinearFeatureAdaptiveExtractionModule$ in Figure 4 (a)).
(a) Feature Acquisition module.
(b) Nonlinear Feature Adaptive Extraction Module.
Figure 4: Temporal-Nonlinear feature extraction module.
Temporal Feature Extraction Module Inspired by the concept of SPP [36], the temporal feature extraction module employs one-dimensional (1D) convolutions with varying dilation factors to capture features and patterns at different resolutions within the input multivariate time series data. This approach equips the model with the capability to discern local feature variations: with small dilation factors, the model convolves within a small receptive field, facilitating more precise identification of local patterns and structures, thus aiding in enhancing the network’s comprehension of complex patterns. Conversely, with large dilation factors, the model employs larger convolutional kernels to capture features within a broader receptive field, thereby facilitating the identification of local contextual relationships within the input sequence.
Additionally, by aggregating feature and pattern information extracted at different resolutions at the same location, it ensures positional invariance of the data during the processing. To encapsulate global trends and patterns, our model incorporates adaptive average pooling operations. This technique is instrumental in distilling the overarching trends and patterns from the time series data, thus enriching the model’s comprehensive grasp of the temporal dynamics presented.
@@ 3.3 Feature acquisition
Nonlinear Feature Adaptive Extraction Module Figure 1 illustrates that dependencies between different time observations in time series data may exhibit nonlinear and non-uniform distributions. To extract nonlinear features from complex time series, we designed a gated deformable convolution (GDC) module. Specifically, we added a gating mechanism to the deformable convolution, which further enhances the model’s flexibility and capability, helping it to better capture nonlinear features in complex time series. By adjusting the positions of the convolutional kernels, deformable convolution enhances the model’s ability to capture local nonlinear features, better handling irregular features and patterns, thereby improving its perception of local nonlinear patterns, as shown in Figure 4 (b). This capability is particularly important for time series data with significant nonlinear and dynamic feature changes.
Specifically, we stack the extracted multi-resolution feature maps and the output maps from adaptive average pooling into a tensor $X_{sd}$, which serves as the input to the nonlinear feature adaptive extraction module. We employ GDC module to learn the nonlinear information between different observations and position offset information in feature graphs with different resolutions, focusing attention on regions with richer information through twist sampling networks.
This operation enables the model to more accurately capture and understand the nonlinear features and relationships between different variables that vary between short-term and long-term in time series, while removing redundant feature information. The mathematical expression of the nonlinear feature extraction module is as follows:
@@
where $DefConv(p_{0})$ is the value of the output feature map at location $p_{0}$. $x(\cdot)$ represents the input feature map. $\mathcal{W}(p_{n})$ denotes the weights of the convolutional kernel. $\triangle p_{n}$ is the learned offset for each $p_{n}$, differing for each point, allowing the convolutional kernel to adaptively adjust its sampling positions on the input feature map. For more detailed theoretical explanations regarding deformable convolutions, please refer to reference [17]. The nonlinear feature adaptive extraction is described by the following equations:
$\begin{split}&X_{sd}=Stack(\chi_{dilated},\chi_{avg\_pool}),\\ &Gate=Sigmoid(Conv2D(X_{sd})),\\ &X_{dc}=Gate*DefConv(X_{sd}),\end{split}$ (10)
where $Conv2D$ denotes a two-dimensional convolution operation, while $sigmoid$ represents the activation function.